2. Compute the following integrals: OO 1 da δ(-1) 1 2 (a) dx 8 (x2) sin(a) (b) dx 8(x -5) nx (c) 88

please help and ill give good rating
Answer #1

Fen) = 341 = (x2-1)(x+1) = (3-1) (n+1) (4+1) (91-)) f(n=0, 121, 1,2-1, n = -) e Hy = fi fen) = 443_0 = 443 f(211) = 4 f(22)please be careful about part b .plesse check answer some other source and let me know .thank you